# Mipha: A Comprehensive Overhaul of Multimodal Assistant with Small Language Models

### 13 Mar 2024 v2

@@ 1 Introduction
Intuitively, reducing the computational demands of the language model could lead to a significant decrease in overall inference costs. The rapid advancements by the open-source community in developing Small Language Models (SLMs), with parameter sizes ranging from 1 billion to 3 billion (i.e., Phi Gunasekar et al. (2023); Li et al. (2023e); Javaheripi et al. (2023)), enable us to construct vision-language models more compact than traditional MLLMs. However, as the parameter count of these smaller language models is reduced, their capabilities often do not measure up to those of their larger counterparts, potentially diminishing the effectiveness of small multimodal models. This motivation drives us to explore the development of multimodal assistants using small language models that can deliver competitive performance comparable to their counterparts Dai et al. (2023b); Bai et al. (2023); Ye et al. (2023a, b); Liu et al. (2023a), which rely on large language models.
@@ 4.1.1 Which Small Language Model should we adopt?
As the "brain" of multimodal models, the selection of pretrained small language models is crucial for crafting high-performing Multimodal Small Language Models (MSLMs). We evaluate a spectrum of recently introduced Small Language Models (SLMs) with sizes ranging from 1.3B to 2.8B parameters. This includes models such as Phi-2 (2.7B) Javaheripi et al. (2023), Phi-1.5 (1.3B) Li et al. (2023e), Qwen-1.5 (1.8B) Bai et al. (2023), Pythia (available in both 1.4B and 2.8B sizes) Biderman et al. (2023), and StableLM (1.6B) Stability (2024). For analytical clarity, we classify these SLMs into two categories: those models possess over 2.7B parameters, and the remainder, feature less than 2B parameters.
@@ 4.3 Optimization
Optimization strategies play a pivotal role in shaping the generalizability of a model. Most MLLMs Liu et al. (2023b, a); Zhu et al. (2024a); Ye et al. (2023b) typically adopt a two-stage optimization pipeline. The first stage, or the pretraining stage, is dedicated to harmonizing vision and language features, establishing a solid foundation for multimodal understanding. The subsequent stage focuses on instructing the model to follow users’ instructions. This section delves into the training paradigms of MLLMs, exploring the necessity of finetuning the language model and visual backbones. Additionally, we evaluate the feasibility and effectiveness of parameter-efficient fine-tuning methods, such as LoRA, in this context.
@@
Qwen-VL Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
Qwen-VL-Chat Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
@@ 5.2 Comparison with SOTA Models
Qwen-VL Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
Qwen-VL-Chat Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -

### 15 Mar 2024 v3

@@ 1 Introduction
Intuitively, reducing the computational demands of the language model could lead to a significant decrease in overall inference costs. The rapid advancements by the open-source community in developing Small Language Models (SLMs), with parameter sizes ranging from 1 billion to 3 billion (i.e., Phi Gunasekar et al. (2023); Li et al. (2023e); Javaheripi et al. (2023)), enable us to construct vision-language models more compact than traditional MLLMs. However, as the parameter count of these smaller language models is reduced, their capabilities often do not measure up to those of their larger counterparts, potentially diminishing the effectiveness of small multimodal models. This motivation drives us to explore the development of multimodal assistants using small language models that can deliver competitive performance comparable to their counterparts Dai et al. (2023b); Bai et al. (2023b); Ye et al. (2023a, b); Liu et al. (2023a), which rely on large language models.
@@ 4.1.1 Which Small Language Model should we adopt?
As the "brain" of multimodal models, the selection of pretrained small language models is crucial for crafting high-performing Multimodal Small Language Models (MSLMs). We evaluate a spectrum of recently introduced Small Language Models (SLMs) with sizes ranging from 1.3B to 2.8B parameters. This includes models such as Phi-2 (2.7B) Javaheripi et al. (2023), Phi-1.5 (1.3B) Li et al. (2023e), Qwen-1.5 (1.8B) Bai et al. (2023b), Pythia (available in both 1.4B and 2.8B sizes) Biderman et al. (2023), and StableLM (1.6B) Stability (2024). For analytical clarity, we classify these SLMs into two categories: those models possess over 2.7B parameters, and the remainder, feature less than 2B parameters.
@@ 4.3 Optimization
Optimization strategies play a pivotal role in shaping the generalizability of a model. Most MLLMs Liu et al. (2023b, a); Zhu et al. (2024a); Ye et al. (2023b) typically adopt a two-stage optimization pipeline. The first stage, or the pretraining stage, is dedicated to harmonizing vision and language features, establishing a solid foundation for multimodal understanding.
Table 2: Multi-modal evaluation on 9 benchmarks: VQAv2 Goyal et al. (2017); GQA Hudson and Manning (2019); $\text{SQA}^{\text{I}}$: ScienceQA-IMG Lu et al. (2022); $\text{VQA}^{\text{T}}$: TextVQA Singh et al. (2019); MME-P: MME perception Fu et al. (2023); MME-C: MME cognition Fu et al. (2023); MMB: MMBench Liu et al. (2023d); MM-Vet Yu et al. (2023); POPE Li et al. (2023d). $*{}^{*}$he training images/annotations of the datasets are observed during training. ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Column Res. is the image resolution of visual backbone. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
[…]
Multimodal Large Language Models
BLIP-2 Li et al. (2023c) Vicuna (13B) 224 129M - 65.0 41.0 61.0 42.5 1293.8 290.0 - 22.4 85.3
InstructBLIP Dai et al. (2023b) Vicuna (7B) 224 129M 1.2M - 49.2 60.5 50.1 - - 36 26.2 -
InstructBLIP Dai et al. (2023b) Vicuna (13B) 224 129M 1.2M - 49.5 63.1 50.7 1212.8 291.8 - 25.6 78.9
Shikra Chen et al. (2023b) Vicuna (13B) 224 600K 5.5M $*{}^{*}$ 58.8 - -
IDEFICS-9B Laurençon et al. (2023) LLaMA (7B) 224 353M 1M 50.9 38.4 - 25.9 - - 48.2 - -
IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 - 30.9 - - 54.5 - -
Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 67.1 $*{}^{*}$ 38.2 - -
Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 68.2 $*{}^{*}$ 1487.5 360.71 60.6 - -
mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 68.7 58.2 1450.2 313.2 64.5 36.2 85.8
LLaVA-1.5 Liu et al. (2023a) Vicuna (7B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 66.8 58.2 1510.7 316.1 64.3 30.5 85.9
LLaVA-1.5 Liu et al. (2023a) Vicuna (13B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 71.6 61.3 1531.3 295.4 67.7 35.4 85.9
Multimodal Small Language Models
MobileVLM-1.7B Chu et al. (2023) M-LLaMA (1.4B) 336 558K 665K - 56.1 57.3 41.5 1196.2 - 53.2 - 84.5
MobileVLM-3B Chu et al. (2023) M-LLaMA (2.7B) 336 558K 665K - 59.0 61.2 47.5 1288.9 - 59.6 - 84.9
MobileVLM-v2-1.7B Chu et al. (2024) M-LLaMA (1.4B) 336 1.2M 2.4M - 59.3 66.7 52.1 1302.8 - 57.7 - 84.3
MobileVLM-v2-3B Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 70.0 57.5 1440.5 - 63.2 - 84.7
MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 70.3 57.0 1431.3 - 68.0 35.9 85.7
Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 70.9 - 1488.8 289.3 68.6 - 86.6
Mipha-1.6B Phi-1.5 (1.3B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 58.3 45.6 1203.1 247.9 57.7 23.5 86.9
Mipha-2.4B Gemma (2.0B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 65.3 52.4 1397.1 265.7 59.4 29.9 86.6
Mipha-3B Phi-2 (2.7B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 70.9 56.6 1488.9 295.0 69.7 32.1 86.7
The subsequent stage focuses on instructing the model to follow users’ instructions. This section delves into the training paradigms of MLLMs, exploring the necessity of finetuning the language model and visual backbones. Additionally, we evaluate the feasibility and effectiveness of parameter-efficient fine-tuning methods, such as LoRA, in this context.
@@
Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
@@ 5.2 Comparison with SOTA Models
Qwen-VL Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
Qwen-VL-Chat Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
@@ 6 Conclusion
Appendix A Limitation
While this work has extensively and thoroughly explored the design space of Multimodal Small Language Models (MSLMs), our analysis predominantly draws upon methodologies from current models, such as Phi-1.5/2 Li et al. [2023e], Javaheripi et al. [2023], StableLM Stability [2024], Pythia Biderman et al. [2023], and Gemma Google [2024], which are considered smaller language models. Given the rapid evolution of the field, the findings and conclusions of our study might not remain applicable to newer language models that emerge. Moreover, we recognize significant potential for further enhancements in Mipha, particularly in areas like Optical Character Recognition (OCR) and mathematical problem-solving. We plan to tackle these shortcomings in future research endeavors.
Appendix B Negative Social Impact
Despite Mipha showcasing advanced multimodal comprehension capabilities, it is not immune to the negative societal impacts observed in existing multimodal models. Foremost, Mipha has the potential to produce unreliable responses and may inadvertently perpetuate societal biases, including gender biases. Additionally, there exists the risk of Mipha being manipulated by users to generate harmful content.
Appendix C Implementation Details
The training process and dataset adopted by Mipha are substantially identical to LLaVA-1.5 Liu et al. [2023a]. Specifically, Mipha adopts a two-stage training pipeline, i.e., the pretraining stage and the instruction tuning stage. Initially, we align the vision features from SigLIP Zhai et al. [2023] with the text embedding from Small Language Models (SLMs) Li et al. [2023e], Javaheripi et al. [2023]. Subsequently, we apply visual instruction tuning to fully harness the capabilities of the Multimodal Small Language Models (MSLMs) across various multimodal tasks. We adopt the same cross-entropy loss during both stages for the next token prediction. In the pretraining stage, only the parameters of projector (a two-layer MLP) are updated on LCS-558K Liu et al. [2023a]. The learning rate of 1e-3 and a batch size of 256 are adopted at this stage. In the instruction tuning stage, we train all parameters of Mipha for 2 epochs on a collection of mixture datasets with 665K samples, which is organized by Liu et al. [2023a]. At this stage, we set the learning rate as 2e-5 and the batch size as 128. For both stages, we utilize the AdamW optimizer with no weight decay, characterized by momentum parameters of 0.9 and 0.98, and an epsilon value of 1e-7. We provide overall training settings in Fig 5. Additionally, for our model variants, we employ the same hyperparameters as Mipha, except for setting the epoch to 1 during the finetuning stage. We employ publicly available datasets to conduct instruct tuning for LMs. Specifically, we utilize the Alpaca dataset Taori et al. [2023] to supervised fine-tune Phi-2. For RHLF Ouyang et al. [2022] and Safe-RLHF Dai et al. [2023a], we employ the PKU-SafeRLHF-10K dataset Dai et al. [2023a] to finetune the models.
Table 5: Hyper-parameters for training Mipha.
Configuration Pretraining Instruction Tuning
ViT init. SigLIP 384px/14 SigLIP 384px/14
Projector init. Random Pretrain stage
Image resolution $384\times 384$
ViT sequence length 729
LLM sequence length 2048
Optimizer AdamW
Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1e^{-7}$
Peak learning rate $1e^{-4}$ $2e^{-5}$
Training Epochs 1 2
Global batch size 256 128
Learning rate schedule Cosine
Weight decay 0
Warm-up ratio 0.03
Numerical precision $\mathtt{bfloat16}$
Activation checkpointing
Figure 6: Additional Experimental Results: Selection of Small Language Models (Left) & Scaling Image Resolution (Right).
C.1 More Experiments
C.1.1 Selection of Small Language Models.
Beyond the experiments presented in Fig.  2 (left) of the main text, we have expanded our evaluation to include the performance of the Gemma-2B model Google [2024]. We compared Gemma-2B with several smaller language models: Qwen-1.8B Bai et al. [2023a], Pythia-1.4B Biderman et al. [2023], StableLM-1.6B Stability [2024], and Phi-1.5-1.3B Li et al. [2023e]. The results of these comparisons are illustrated in Fig 6 (left). The findings suggest that multimodal models leveraging Gemma-2B are versatile, achieving top-tier performance across the majority of benchmarks when compared to alternative methods. As one of the most recent additions to the small language model category, its performance on language-based tasks has exceeded that of LLaMA-2 Touvron et al. [2023b]. We posit that Gemma-2B’s robust performance makes it an excellent foundational model for the development of MSLMs.
C.1.2 Scaling Image Resolution.
In addition to CLIP Radford et al. [2021], we also conduct experiments using SigLIP Zhai et al. [2023] with resolutions of 384px and 448px. Our findings, as illustrated inFig 6 (right), reveal that the SigLIP model with a 384px resolution outperforms its 448px counterpart in five out of eight benchmarks Hudson and Manning [2019], Singh et al. [2019], Goyal et al. [2017], Lu et al. [2022], Li et al. [2023d], Yu et al. [2023], Liu et al. [2023d], Fu et al. [2023]. This suggests that simply increasing image resolution does not consistently improve the model’s performance across a diverse set of multimodal tasks. Additionally, with the exception of MM-Vet Yu et al. [2023], the two SigLIP variants show minimal variance in performance on the other seven benchmarks. This observation reinforces our assertion that higher image resolutions do not necessarily equate to better results in MSLMs.
C.2 Qualitative Examples
In Fig 7, 8 and 9, we present several cases from various perspectives such as commonsense reasoning, OCR, image description, object recognition, joke comprehension, and multilingual understanding. Mipha-3B demonstrates powerful detailed visual understanding by comprehensively answering questions based on the information in the images.
Figure 7: Some cases of Commonsense Reasoning & OCR by Mipha-3B. For Commonsense Reasoning, Mipha-3B successfully identifies the Great Wall of Chain and the Great Pyramid of Giza in the image, going further to provide more detailed descriptions of them. Additionally, Mipha-3B also demonstrates strong OCR capabilities, such as successfully recognizing “since 1994” in the bottom right subfigure. RED means the critical part of the answer.
Figure 8: Some cases of Image Description & Object Recognition by Mipha-3B. Mipha-3B is capable of inferring potential occurrences based on the scene, such as “a drag race event”, showcasing its robust visual perception abilities. Additionally, Mipha-3B demonstrates strong object recognition capabilities, including aspects such as counting, existence, and orientation. RED means the critical part of the answer.
Figure 9: Some cases of Jokes Comprehension & Multilingual Understanding by Mipha-3B. Mipha-3B comprehends intriguing elements within the image and demonstrates multilingual abilities across Chinese and French. RED means the critical part of the answer.